OpenStudy (anonymous):

For the simple harmonic motion equation d=2sin((pi/3)t), what is the frequency? If necessary, use the slash (/) to denote a fraction.

OpenStudy (kropot72):

The generalised form of this equation is\[d=2\sin 2\pi ft\] where 2 * pi * f is the angular frequency \[2\pi f=\frac{\pi}{3}\] \[f=\frac{\pi}{3}\times \frac{1}{2\pi}=\frac{1}{6}Hz\]